# Low-Light Video Enhancement via Spatial-Temporal Consistent Decomposition

### v2

@@ 1 Introduction
Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions (Xu et al., 2022; Wang et al., 2021), improving their visual quality while reducing noise. This technique can be applied in wide-ranging applications, such as portrait photography on mobile devices (Ignatov et al., 2017; Hasinoff et al., 2016), nighttime face recognition (Ma et al., 2022; Wang et al., 2022) and vehicle detection (Fu et al., 2023a; Wu et al., 2022). The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally. Moreover, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
@@
One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data. According to the linear Lambertian model for intrinsic image decomposition (Yuille, 2012; Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be expressed as the element-wise product ($\otimes$) of its albedo component and shading component. The albedo part is normally view-independent, capturing the intrinsic appearance, whereas the shading component is commonly view-dependent, influenced by lighting conditions and the surface normal’s direction. Inspired by the Lambertian model, we propose that a frame in the LLVE task can similarly be decomposed as $\boldsymbol{I}=\boldsymbol{L}\otimes\boldsymbol{R}$, where $\boldsymbol{L}$ represents the view-dependent term, and $\boldsymbol{R}$ corresponds to the view-independent component.
@@
Additionally, we introduce a dual-network structure to enforce the consistency of feature representations across frames, facilitating the consistent synthesis of decomposition in turn. Two frames are used as inputs, and their features are propagated within the designed Cross-Frame Interaction Module (CFIM). CFIM differs from other similar architectures in the restoration task (Lv et al., 2023) in terms of the fusion manner. CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives. Moreover, CFIM selectively incorporates knowledge from randomly chosen neighboring frames into the backbone (i.e., the part without feature propagation) during training. In this way, CFIM help the backbone in learning how to adaptively utilize knowledge from various cross-frame features, resulting in a more robust backbone.
@@ 2 Related Work
Low-Light Image Enhancement. To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis. In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques (Jiang et al., 2021; Yang et al., 2021) have made significant advancements, with a primary focus on supervised approaches.
Low-Light Video Enhancement. In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment. Various approaches have been proposed in this context (Chen et al., 2019; Triantafyllidou et al., 2020; Ye et al., 2023; Lv et al., 2023; Xu et al., 2023a; Fu et al., 2023a). Liu et.al. (Liu et al., 2023) and Liang et.al. (Liang et al., 2023) used prior event information to learn enhancement mapping for brightening videos. Xu et.al. (Xu et al., 2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos. Recently, Fu et.al. (Fu et al., 2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it. However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
Several datasets for video enhancement, encompassing both static (Chen et al., 2019; Jiang and Zheng, 2019; Wang et al., 2019; Triantafyllidou et al., 2020) and dynamic motions (Wang et al., 2021; Fu et al., 2023a), have been introduced. In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets. Our decomposition method explicitly and consistently models view-dependent and -independent for all frames.
@@ 3.1 Decomposition Model
Motivation. According to the linear Lambertian model (Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be formulated as the element-wise product of its albedo and shading component. Here, the albedo part is view-independent, capturing the intrinsic appearance, while the shading component is view-dependent. Inspired by this, when presented with an image $\boldsymbol{I}$ (in this paper, we denote the low-light image as $\boldsymbol{I}_{d}$ and the normal-light image as $\boldsymbol{I}_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into view-dependent part $\boldsymbol{L}$ and view-independent part $\boldsymbol{R}$, as
@@
where $\mathcal{O}(\boldsymbol{L})$ and $\mathcal{O}(\boldsymbol{R})$ denote the constraints for $\boldsymbol{L}$ and $\boldsymbol{R}$ (we will introduce our designed new constraints in Sec. 3.2). When applying this model to sequential video data, the objective related to the decomposition of video data can be summarized as
@@
where $\mathbb{E}$ is the average operation, $T$ is the number of frame, and $\boldsymbol{L}_{t}$/$\boldsymbol{R}_{t}$ is the decomposition output at the time index of $t$. Compared with Equation 2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
@@ 3.2 Spatial-Temporal Consistent Decomposition
Motivation. Video can be conceptualized as static/dynamic 3D multi-view data (Wang et al., 2023b, a). Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations (Wang et al., 2023a). Thus, once we establish correspondence relationships among $\boldsymbol{I}_{d,t}$ for all $t$, we can apply the view-independent constraint to derive $\hat{\boldsymbol{R}}_{n,t}$, as it represents the intrinsic properties of the target scene. Subsequently, obtaining $\hat{\boldsymbol{L}}_{n,t}$ becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
@@
Given two frames $\boldsymbol{I}_{n,t_{1}}$ and $\boldsymbol{I}_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ (Edstedt et al., 2023). We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$. Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $\boldsymbol{I}_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $\boldsymbol{I}_{n,t_{2}}$. Furthermore, each correspondence is associated with an uncertainty value $u_{m}$, determined through a data-driven approach.
@@ 3.3 Dual Network for Consistent Decomposition
Overview and Motivation. To enhance the consistency of decomposition, it’s crucial for various frames within a video to mutually share features. By incorporating shared features and simultaneous supervision, the features used for synthesizing decomposition parameters across different frames can exhibit greater consistency. When the video is enhanced frame by frame without feature propagation, the decomposition results tend to be suboptimal due to the inherently challenging nature of maintaining consistency. This observation is empirically validated in our ablation study. Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation (Wang et al., 2021; Fu et al., 2023a). However, these strategies come with a notable cost in terms of aligning features, potentially affecting inference efficiency.
@@
Feature Propagation via Spatial-varying Fusion. Suppose the feature of $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ are denoted as $\boldsymbol{f}_{d,t_{1}}$ and $\boldsymbol{f}_{d,t_{2}}$ which is extracted from the same encoder $M_{E}$ in the network $f$. To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space. A traditional attention operation (Vaswani et al., 2017) typically consists of the query vector $\boldsymbol{Q}$, the key vector $\boldsymbol{K}$, and the value vector $\boldsymbol{V}$. The attention relationship is established using $A(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top})\times\boldsymbol{V}$, where $\times$ represents matrix multiplication. To enable cross-frame attention, we process the feature via dual paths, as
@@
The Role of the Cross-attention in CFIM. The cross-attention mechanism facilitates information propagation across different frames, aligning with our training procedure that requires spatial consistency for $\boldsymbol{R}$. Moreover, during training, CFIM brings varying filtered features from randomly selected reference images ($\boldsymbol{I}_{d,t_{2}}$) into the backbone (i.e., the part without feature propagation), where the input is the image at the current time step $\boldsymbol{I}_{d,t_{1}}$. This setup enables CFIM to guide the backbone in learning how to adaptively leverage diverse knowledge from reference images, resulting in robust and effective feature spaces within the backbone, without the need for expensive temporal alignment. In other words, the backbone performs strongly without a reference image during inference (though, of course, having a suitable reference image to provide additional useful information generally improves performance), as confirmed by the ablation study. Additionally, CFIM can leverage useful cross-frame information when it complements the features in the backbone, further enhancing the decomposition process.
@@ 4.1 Datasets
Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations, including SMID (Chen et al., 2019), SDSD (Wang et al., 2021), DID (Fu et al., 2023a), and DAVIS (Pont-Tuset et al., 2017).
@@ 4.2 Implementation Details
Experimental Details. We conducted experiments on all datasets using the same network structure. $T$ is set as 5 in the experiment. All modules were trained end-to-end, with the learning rate initialized at $4e^{-4}$ for all layers, adapted by a cosine learning scheduler. The batch size used was 4. Correspondences were computed using the SOTA method DKM (Edstedt et al., 2023). We used the pre-trained weights of DKM during training because it is designed for general indoor and outdoor scenes. Its generalization ability is supported by its extensive training data and state-of-the-art training strategy, as demonstrated in the original DKM paper and subsequent studies (Zhu and Liu, 2023; Edstedt et al., 2024) (through evaluations on unseen real-world indoor and outdoor scenes).
@@ 4.3 Comparison and Baselines
We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN (Lv et al., 2018), SMID (Chen et al., 2019), SMOID (Jiang and Zheng, 2019), SDSDNet (Wang et al., 2021), DP3DF (Xu et al., 2023a), StableLLVE (Zhang et al., 2021), LLVE-SEG (Liu et al., 2023), LLVE-CFA (Chhirolya et al., 2022), BLLRVE (Zhang et al., 2024), VRGSS (Li et al., 2023), ZRLLE-PQP (Wang et al., 2024), and JDAE (Shi et al., 2024). We also compared our method with SOTA LLIE methods for individual frames, including SNR (Xu et al., 2022), SMG (Xu et al., 2023b), PairLLE (Fu et al., 2023b), RetinexFormer (Cai et al., 2023). All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee convergence). Note that all baselines are trained on our unified data split for a fair comparison. The splits slightly differ from those in baselines’ original papers, so the scores of baselines may vary from the original papers.
@@
Furthermore, Table 2 provides a summary of the comparative results on DAVIS. In comparison to the degradation synthesis strategy as presented in (Zhang et al., 2021), we have extended our approach to include the degradation of a low-light noise term. As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge. As illustrated in Table 2, our method consistently yields the highest PSNR and SSIM values, reaffirming the effectiveness of our approach.
@@
Evaluation for Temporal Consistency. For video processing, the performance of temporal consistency and stability should be evaluated. Thus, we employ the short-term and long-term temporal loss proposed in (Lai et al., 2018) for such temporal evaluation on different datasets. The wrapping operations among frames are computed on the frames with the normal light. Moreover, the long-term loss is computed every 10 frames. The results are shown in Table 3 (we normalize the frame values into [0, 1]). Obviously, our results have lower temporal loss than the baselines, e.g., higher temporal consistency and stability. Specifically, we assess the temporal consistency of the view-independent term. The results presented in Table 3 further validate the constancy.
@@ 4.5 User Study
To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution. Following most of the existing low-light enhancement works (Xu et al., 2023b; Wang et al., 2023a), we employ the AB-test for the user study. Participants should indicate their preference or select the same option. They should make the decision according to the natural brightness, contrast, and color of each frame; the rich details, fewer artifacts, and the temporal consistency in the video.

### v3

@@ 1 Introduction
Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions Xu et al. (2022); Wang et al. (2021), improving their visual quality while reducing noise. This technique can be applied in wide-ranging applications, such as portrait photography on mobile devices Ignatov et al. (2017); Hasinoff et al. (2016), nighttime face recognition Ma et al. (2022); Wang et al. (2022) and vehicle detection Fu et al. (2023a); Wu et al. (2022). The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally. Moreover, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
@@
One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data. According to the linear Lambertian model for intrinsic image decomposition Yuille (2012); Narihira et al. (2015), an observed image $\boldsymbol{I}$ can be expressed as the element-wise product ($\otimes$) of its albedo component and shading component. The albedo part is normally view-independent, capturing the intrinsic appearance, whereas the shading component is commonly view-dependent, influenced by lighting conditions and the surface normal’s direction. Inspired by the Lambertian model, we propose that a frame in the LLVE task can similarly be decomposed as $\boldsymbol{I}=\boldsymbol{L}\otimes\boldsymbol{R}$, where $\boldsymbol{L}$ represents the view-dependent term, and $\boldsymbol{R}$ corresponds to the view-independent component.
@@
Additionally, we introduce a dual-network structure to enforce the consistency of feature representations across frames, facilitating the consistent synthesis of decomposition in turn. Two frames are used as inputs, and their features are propagated within the designed Cross-Frame Interaction Module (CFIM). CFIM differs from other similar architectures in the restoration task Lv et al. (2023) in terms of the fusion manner. CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives. Moreover, CFIM selectively incorporates knowledge from randomly chosen neighboring frames into the backbone (i.e., the part without feature propagation) during training. In this way, CFIM help the backbone in learning how to adaptively utilize knowledge from various cross-frame features, resulting in a more robust backbone.
@@ 2 Related Work
Low-Light Image Enhancement. To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis. In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques Jiang et al. (2021); Yang et al. (2021) have made significant advancements, with a primary focus on supervised approaches.
Low-Light Video Enhancement. In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment. Various approaches have been proposed in this context Chen et al. (2019); Triantafyllidou et al. (2020); Ye et al. (2023); Lv et al. (2023); Xu et al. (2023a); Fu et al. (2023a). Liu et.al. Liu et al. (2023) and Liang et.al. Liang et al. (2023) used prior event information to learn enhancement mapping for brightening videos. Xu et.al. Xu et al. (2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos. Recently, Fu et.al. Fu et al. (2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it. However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
Several datasets for video enhancement, encompassing both static Chen et al. (2019); Jiang and Zheng (2019); Wang et al. (2019); Triantafyllidou et al. (2020) and dynamic motions Wang et al. (2021); Fu et al. (2023a), have been introduced. In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets. Our decomposition method explicitly and consistently models view-dependent and -independent for all frames.
@@ 3.1 Decomposition Model
Motivation. According to the linear Lambertian model Narihira et al. (2015), an observed image $\boldsymbol{I}$ can be formulated as the element-wise product of its albedo and shading component. Here, the albedo part is view-independent, capturing the intrinsic appearance, while the shading component is view-dependent. Inspired by this, when presented with an image $\boldsymbol{I}$ (in this paper, we denote the low-light image as $\boldsymbol{I}_{d}$ and the normal-light image as $\boldsymbol{I}_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into view-dependent part $\boldsymbol{L}$ and view-independent part $\boldsymbol{R}$, as
@@
where $\mathcal{O}(\boldsymbol{L})$ and $\mathcal{O}(\boldsymbol{R})$ denote the constraints for $\boldsymbol{L}$ and $\boldsymbol{R}$ (we will introduce our designed new constraints in Sec. 3.2). When applying this model to sequential video data, the objective related to the decomposition of video can be written as
@@
where $\mathbb{E}$ is the average operation, $T$ is the number of frame, and $\boldsymbol{L}_{t}$/$\boldsymbol{R}_{t}$ is the decomposition output at the time index of $t$. Compared with Equation 2, we need to simultaneously guarantee the decomposition quality and the temporal consistency.
@@ 3.2 Spatial-Temporal Consistent Decomposition
Motivation. Video can be conceptualized as static/dynamic 3D multi-view data Wang et al. (2023b, a). Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations Wang et al. (2023a). Thus, once we establish correspondence relationships among $\boldsymbol{I}_{d,t}$ for all $t$, we can apply the view-independent constraint to derive $\hat{\boldsymbol{R}}_{n,t}$, as it represents the intrinsic properties of the target scene. Subsequently, obtaining $\hat{\boldsymbol{L}}_{n,t}$ becomes achievable through utilizing the inherent reconstruction loss, in conjunction with adherence to the continuity assumption.
@@
Given two frames $\boldsymbol{I}_{n,t_{1}}$ and $\boldsymbol{I}_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ Edstedt et al. (2023). We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$. Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $\boldsymbol{I}_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $\boldsymbol{I}_{n,t_{2}}$. Furthermore, each correspondence is associated with an uncertainty value $u_{m}$, determined through a data-driven approach.
@@ 3.3 Dual Network for Consistent Decomposition
Overview and Motivation. To enhance the consistency of decomposition, it’s crucial for various frames within a video to mutually share features. By incorporating shared features and simultaneous supervision, the features used for synthesizing decomposition parameters across different frames can exhibit greater consistency. When the video is enhanced frame by frame without feature propagation, the decomposition results tend to be suboptimal due to the inherently challenging nature of maintaining consistency. This observation is empirically validated in our ablation study. Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation Wang et al. (2021); Fu et al. (2023a). However, these strategies come with a notable cost in terms of aligning features.
@@
Feature Propagation via Spatial-varying Fusion. Suppose the feature of $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ are denoted as $\boldsymbol{f}_{d,t_{1}}$ and $\boldsymbol{f}_{d,t_{2}}$ which is extracted from the same encoder $M_{E}$ in the network $f$. To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space. A traditional attention operation Vaswani et al. (2017) typically consists of the query vector $\boldsymbol{Q}$, the key vector $\boldsymbol{K}$, and the value vector $\boldsymbol{V}$. The attention relationship is established using $A(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top})\times\boldsymbol{V}$, where $\times$ represents matrix multiplication. To enable cross-frame attention, we process the feature via dual paths, as
@@
The Role of the Cross-attention in CFIM. The cross-attention mechanism facilitates information propagation across different frames, aligning with our training procedure that requires spatial consistency for $\boldsymbol{R}$. Moreover, during training, CFIM brings varying filtered features from randomly selected reference images ($\boldsymbol{I}_{d,t_{2}}$) into the backbone (i.e., the part without feature propagation), where the input is the image at the current time step $\boldsymbol{I}_{d,t_{1}}$. This setup enables CFIM to guide the backbone in learning how to adaptively leverage diverse knowledge from reference images, resulting in robust and effective feature spaces within the backbone, without the need for expensive temporal alignment. CFIM also leverages cross-frame information when it complements backbone.
@@ 4.1 Datasets
Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations, including SMID Chen et al. (2019), SDSD Wang et al. (2021), DID Fu et al. (2023a), and DAVIS Pont-Tuset et al. (2017).
@@ 4.2 Implementation Details
Experimental Details. We conducted experiments on all datasets using the same network structure. $T$ is set as 5 in the experiment. All modules were trained end-to-end, with the learning rate initialized at $4e^{-4}$ for all layers, adapted by a cosine learning scheduler. The batch size used was 4. Correspondences were computed using the SOTA method DKM Edstedt et al. (2023). We used the pre-trained weights of DKM during training because it is designed for general indoor and outdoor scenes. Its generalization ability is supported by its extensive training data and state-of-the-art training strategy, as demonstrated in the original DKM paper and subsequent studies Zhu and Liu (2023); Edstedt et al. (2024) (through evaluations on unseen scenes).
@@ 4.3 Comparison and Baselines
We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN Lv et al. (2018), SMID Chen et al. (2019), SMOID Jiang and Zheng (2019), SDSDNet Wang et al. (2021), DP3DF Xu et al. (2023a), StableLLVE Zhang et al. (2021), LLVE-SEG Liu et al. (2023), LLVE-CFA Chhirolya et al. (2022), BLLRVE Zhang et al. (2024), VRGSS Li et al. (2023), ZRLLE-PQP Wang et al. (2024), and JDAE Shi et al. (2024). We also compared our method with SOTA LLIE methods for individual frames, including SNR Xu et al. (2022), SMG Xu et al. (2023b), PairLLE Fu et al. (2023b), RetinexFormer Cai et al. (2023). All methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee convergence). Note that all baselines are trained on our unified data split for a fair comparison. The splits slightly differ from those in baselines’ original papers, so scores of baselines may vary from original papers.
@@
Furthermore, Table 2 provides a summary of the comparative results on DAVIS. In comparison to the degradation synthesis strategy as presented in Zhang et al. (2021), we have extended our approach to include the degradation of a low-light noise term. As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge. As illustrated in Table 2, our method consistently yields the highest PSNR and SSIM values, reaffirming the effectiveness of our approach.
@@
Evaluation for Temporal Consistency. For video processing, the performance of temporal consistency and stability should be evaluated. Thus, we employ the short-term and long-term temporal loss proposed in Lai et al. (2018) for such temporal evaluation on different datasets. The wrapping operations among frames are computed on the frames with the normal light. Moreover, the long-term loss is computed every 10 frames. The results are shown in Table 3 (we normalize the frame values into [0, 1]). Obviously, our results have lower temporal loss than the baselines, e.g., higher temporal consistency and stability. Specifically, we assess the temporal consistency of the view-independent term. The results presented in Table 3 further validate the constancy.
@@ 4.5 User Study
To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution. Following most of the existing low-light enhancement works Xu et al. (2023b); Wang et al. (2023a), we employ the AB-test for the user study. Participants should indicate their preference or select the same option. They should make the decision according to the natural brightness, contrast, and color of each frame; the rich details, fewer artifacts, and the temporal consistency in the video.